A rectangle has a perimeter of 30 inches and it's dimensions are whole numbers what is the maximum possible area of the rectangle in square units
Answer by Alan3354(69443)

A rectangle has a perimeter of 30 inches and it's [sic] dimensions are whole numbers what is the maximum possible area of the rectangle in square units
---------
P = 2(L + W) = 30
L + W = 15
Max area --> 7*8 = 56 sq inches
=================
PS it's = it is